# Towards an Understanding of Context Utilization in Code Intelligence

### v1

@@ 7.1. Metric
The evaluation metric is crucial for measuring the correctness of the prediction results and assessing the model performance. Table LABEL:tab:Metric summarizes the metrics utilized in all 146 papers and details their respective usage scenarios. We classify metrics based on their function to maintain the table’s clarity and focus. Besides, we exclude some simple evaluation metrics, such as memory usage (Yuan et al., 2020) and time cost (Kamiya, 2015).
@@
Ranking metrics assess models’ effectiveness in recommending solutions for code intelligence tasks. The top@k metric (Zhang et al., 2023c; Bangcharoensap et al., 2012; Yang et al., 2024b) measures the proportion of relevant results among the top k recommendations, while MAP (Saha et al., 2013; Zhou et al., 2012; Wang and Lo, 2014), MAR (Zhang et al., 2023c; Youm et al., 2015), MFR (Zhang et al., 2023c), and MRR (Asaduzzaman et al., 2016; Saha et al., 2013; Lam et al., 2015) offer a more detailed evaluation of the ranking quality.
@@
Classification metrics measure model performance using confusion matrices, which categorize predictions into positive and negative as one dimension and the correctness of the model prediction as another. Accuracy (Zhou et al., 2023; Abhinav et al., 2021; Lam et al., 2015), precision (Nafi et al., 2019, 2018; Li et al., 2024f), recall (Qiu et al., 2024; Liang et al., 2024; Liao et al., 2024) and F1 score (Nguyen et al., 2012; Shahbazi and Fard, 2023; Zou et al., 2020) are four common classification metrics derived from confusion matrices. Matthews correlation coefficient (MCC) (Chen et al., 2023; Qiu et al., 2024) further considers imbalanced datasets and utilizes confusion matrices to summarize classification quality.
@@
Ranking Top@K (Hit@K) This metric measures the ability of a model to include relevant items in the top K recommendations. (Zhang et al., 2023c; Bangcharoensap et al., 2012; Youm et al., 2015; Swe and Oo, 2019; Zhou et al., 2012; Gao et al., 2023; Motwani and Brun, 2023; Youm et al., 2017; Liu et al., 2023c; Yang et al., 2024b; Wen et al., 2016; Zhao et al., 2024) Defect Detection, Code Generation, Program Repair
MAP (@k) Mean Average Precision (MAP) calculates the average precision for each query for each class. MAP@k calculates the same but within the top k results. (Saha et al., 2013; Lam et al., 2015; Swe and Oo, 2019; Zhou et al., 2012; Wang and Lo, 2014; Youm et al., 2017; Yang et al., 2024b; Lu et al., 2022; Wen et al., 2016; Zhao et al., 2024) Defect Detection
MAR Mean Average Rank (MAR) denotes the mean of the average rank of finding all items. (Zhang et al., 2023c; Youm et al., 2015) Defect Detection
MFR Mean First Rank (MFR) is the mean of localizing the first faulty statement’s rank of all faults under multiple faulty scenarios. (Zhang et al., 2023c) Defect Detection
@@
Classification Accuracy Accuracy represents the proportion of correctly classified samples to the total number of samples. (Lin et al., 2022; Nashid et al., 2023; Chen et al., 2021a; Asaduzzaman et al., 2016; Raychev et al., 2016; Agrawal et al., 2024; Li et al., 2023c; Zhou et al., 2023; Abhinav et al., 2021; Lam et al., 2015; Zou et al., 2020; Liu et al., 2013, 2023b, 2022, 2023a; Phan et al., 2024; Qiu et al., 2024; Lyu et al., 2021; Nashid et al., 2024; Liang et al., 2024; Liao et al., 2024; Wang et al., 2024a; Xu et al., 2021; Jiang et al., 2024; Zhang et al., 2024a) Program Repair, Code Completion, Code Generation, Defect Detection
@@
Precision (@k) Precision measures the proportion of correctly predicted positive instances out of all instances predicted as positive. Precision @k calculates the same but within the top k result. (Agashe et al., 2019; Li, 2020; Lin et al., 2022; Wen et al., 2018; Yu et al., 2023; Asaduzzaman et al., 2016; Nguyen et al., 2012; Liu et al., 2013; Zhou et al., 2019; Yuan et al., 2020; Zügner et al., 2021; Bowman and Huang, 2020; Ghosh and Kuttal, 2018; Zhang et al., 2023b; Chen et al., 2023; Liu et al., 2022, 2023a, 2023c; Qiu et al., 2024; Liang et al., 2024; Liao et al., 2024; Wen et al., 2024b; Xu et al., 2021; Lu et al., 2022; Wang et al., 2023; Nafi et al., 2019, 2018; Li et al., 2024f; Ma et al., 2024a; Zhang et al., 2024a; Jimenez et al., 2024) Code Generation, Program Repair, Code Completion, Clone Detection, Defect Detection
@@
Recall (@k) Recall represents the proportion of correctly predicted positive samples among all positive samples. Recall@K measures the same but within the top k suggestions instead of positive samples. (Agashe et al., 2019; Li, 2020; Lin et al., 2022; Yu et al., 2023; Asaduzzaman et al., 2016; Saha et al., 2013; Nguyen et al., 2012; Wang et al., 2009; Zou et al., 2020; Liu et al., 2013; Zhou et al., 2019; Yuan et al., 2020; Zügner et al., 2021; Bowman and Huang, 2020; Ghosh and Kuttal, 2018; Zhang et al., 2023b; Chen et al., 2023; Liu et al., 2022, 2023a, 2023c; Qiu et al., 2024; Liang et al., 2024; Liao et al., 2024; Xie et al., 2024; Li et al., 2024c; Wen et al., 2024b; Bibaev et al., 2022; Xu et al., 2021; Nafi et al., 2019, 2018; Li et al., 2024f; Ma et al., 2024a; Zhang et al., 2024a) Code Generation, Program Repair, Code Completion, Defect Detection, Clone Detection
@@
F1-score F1-score comprehensively considers the model’s ability by measuring harmonic mean of precision and recall. (Liu et al., 2023c; Li, 2020; Zhou et al., 2023; Lin et al., 2022; Yu et al., 2023; Asaduzzaman et al., 2016; Nguyen et al., 2012; Shahbazi and Fard, 2023; Zou et al., 2020; Liu et al., 2013; Zhou et al., 2019; Yuan et al., 2020; Zügner et al., 2021; Bowman and Huang, 2020; Zhang et al., 2023b; Chen et al., 2023; Liu et al., 2022, 2023a; Ding et al., 2023a; Lyu et al., 2021; Liang et al., 2024; Liao et al., 2024; Cheng et al., 2024; Wen et al., 2024b; Xu et al., 2021; Jiang et al., 2024; Nafi et al., 2019, 2018; Li et al., 2024f; Zhang et al., 2024a; Liu et al., 2024a) Code Generation, Code Completion, Code Summarization, Program Repair, Clone Detection, Defect Detection
@@
MCC Matthews correlation coefficient (MCC) evaluates binary classification performance using the confusion matrix. This metric provides a balanced score even when the class sizes are imbalanced. (Chen et al., 2023; Qiu et al., 2024; Wen et al., 2024b; Jiang et al., 2024) Defect Detection
@@
Similarity BLEU BLEU computes the similarity between prediction and ground truth sequences by accessing the n-gram overlap. (Agashe et al., 2019; Nashid et al., 2023; Yu et al., 2020; Zhou et al., 2023; Washio and Miyao, 2022; Bansal et al., 2023b; Shahbazi and Fard, 2023; Xu et al., 2022; Nie et al., 2021; Wang et al., 2021b; Vu et al., 2023; Shahbazi et al., 2021; Guo et al., 2023a; Ding et al., 2023b; Haque et al., 2020; Wang et al., 2021a; Bansal et al., 2021, 2023c, 2023a; Matmti and Fard, 2023; Lin et al., 2021; Gao et al., 2023; Liu et al., 2021; Zhang et al., 2021; Lyu et al., 2021; Wang et al., 2024g, f; Mandli et al., 2024; Guo et al., 2024a; Wang et al., 2023; Li et al., 2024a; Eliseeva et al., 2023; Imani et al., 2024; Zhang et al., 2024h; Abdu et al., 2024; Zhou et al., 2024) Code Generation, Code Summarization, Program Repair, Commit Message Generation, Code Completion
@@
CodeBLEU CodeBLEU is a bleu-modified metric tailored for code-related tasks. CodeBLEU evaluates the similarity of code by weighting the combination of n-gram match, data-flow match, and AST match. (Liu et al., 2023c; Phan et al., 2024; Wang et al., 2024g; Tan et al., 2024; Li et al., 2024d) Code Generation, Code Completion
@@
EM Exact Match (EM) measures the proportion of predictions that match the ground truth exactly. (Li et al., 2023c; Zhou et al., 2023; Pei et al., 2023; Ding et al., 2023b; Shrivastava et al., 2023b; Zhang et al., 2023a; Shrivastava et al., 2023a; Guo et al., 2023b; Ding et al., 2023a; Liu et al., 2023c; Agashe et al., 2019; Phan et al., 2024; Wu et al., 2024b; Li et al., 2024e; Nashid et al., 2024; Liang et al., 2024; Wang et al., 2024g; Cheng et al., 2024; Tan et al., 2024; Li et al., 2024d; Wang et al., 2024d; Liu et al., 2024b; Deng et al., 2024; Lu et al., 2022; Eliseeva et al., 2023; Prenner and Robbes, 2024; Liu et al., 2024a) Code Generation, Code Completion, Commit Message Generation
@@
ES Edit Similarity (ES) measures the similarity between two strings, defined as the minimum number of single-character edits. (Pei et al., 2023; Shrivastava et al., 2023b; Zhang et al., 2023a; Guo et al., 2023b; Ding et al., 2023a; Wu et al., 2024b; Li et al., 2024e; Liang et al., 2024; Wang et al., 2024g; Cheng et al., 2024; Tan et al., 2024; Wang et al., 2024d; Liu et al., 2024b; Deng et al., 2024; Lu et al., 2022; Eliseeva et al., 2023; Wang et al., 2022; Liu et al., 2024a) Code Generation, Code Completion, Commit Message Generation
SM Syntax Match (SM) quantifies the match of subtrees within the code.  (Li et al., 2024d) Code Completion
@@
ROUGE ROUGE utilizes the longest common subsequence to measure the matching degree between the generated and reference results. (Yu et al., 2020; Xu et al., 2022; Nie et al., 2021; Wang et al., 2021b; Vu et al., 2023; Shahbazi et al., 2021; Zhou et al., 2019; Guo et al., 2023a; Haque et al., 2020; Wang et al., 2021a; Bansal et al., 2021, 2023a; Matmti and Fard, 2023; Lin et al., 2021; Gao et al., 2023; Liu et al., 2021; Zhang et al., 2021; Lyu et al., 2021; Wang et al., 2024f; Mandli et al., 2024; Wang et al., 2023; Li et al., 2024a; Imani et al., 2024; Abdu et al., 2024; Zhou et al., 2024) Code Generation, Code Summarization, Commit Message Generation, Program Repair
@@
METEOR METEOR computes the harmonic meaning between unigram precision and recall. (Bansal et al., 2023b; Shahbazi and Fard, 2023; Xu et al., 2022; Nie et al., 2021; Wang et al., 2021b; Vu et al., 2023; Shahbazi et al., 2021; Zhou et al., 2019; Guo et al., 2023a; Wang et al., 2021a; Bansal et al., 2023c, a; Matmti and Fard, 2023; Gao et al., 2023; Liu et al., 2021; Zhang et al., 2021; Wang et al., 2024f; Mandli et al., 2024; Guo et al., 2024a; Li et al., 2024a; Zhang et al., 2024h; Abdu et al., 2024; Zhou et al., 2024; Zhang et al., 2024a) Code Summarization, Commit Message Generation, Program Repair
@@
CIDEr CIDEr is a consensus-based evaluation metric used in image captioning tasks to measure whether the model captures important messages. (Wang et al., 2021a; Lyu et al., 2021) Code Generation, Code Summarization
BM25 score Best Matching 25 (BM25) score estimates the relevance of texts based on term frequency, inverse document frequency, and document length normalization. (Zhang et al., 2023d) Code Generation
RIBES RIBES estimates models based on rank correlation and evaluates the distance between the prediction and reference. (Lyu et al., 2021) Code Generation
Model-Related Perplexity Perplexity measures how well a probability distribution predicts a sample. This metric computes the exponentiation of the average negative log-likelihood of the predicted probabilities. (Lutellier et al., 2020; Chen et al., 2021a; Lu et al., 2022) Program Repair, Code Completion
RImp Relative Improvement (RImp) calculates the improvement of solving tasks between the old and new techniques. (Zhang et al., 2023c; Li et al., 2024b) Defect Detection, Program Repair
@@
Compiler-Based Pass@k This metric calculates the proportion of solved questions that pass the unit test. (Zan et al., 2023, 2022a; Tang et al., 2023; Zhou et al., 2023; Zhang et al., 2023d; Bi et al., 2024b; Yu et al., 2024; Xie et al., 2024; Li et al., 2024c; Wang et al., 2024g; Zhang et al., 2024b; Zheng et al., 2024c; Wu et al., 2024c; Guan and Treude, 2024; Zhuo et al., 2024) Code Generation, Code Completion
@@
ValRate Static Validity Rate (ValRate) calculates the performance of models in reducing dependency errors by evaluating the proportion of generated code that successfully passes a static check for dependency errors. (Wang et al., 2024g) Code Generation
@@
ACR Average Coverage Ratio (ACR) calculates the average ratio of successfully predicted statements in all buggy statements. (Yang et al., 2024b) Defect Detection
@@
Dependency Coverage Dependency Coverage means the proportion of dependencies that appear in ground-truth functions while covered by the prediction. (Wang et al., 2024g) Code Generation
API Coverage API Coverage reports the ratio of API coverage in code snippets. (Zhang et al., 2024f) Defect Detection
Task-Specific Full Repair Full Repair evaluates the ability to fix a broken program, including the location of an error statement and the fix of this statement. (Li et al., 2023c) Program Repair
@@
Similarity metrics assess how closely the model’s predictions match the ground truth. Bleu (Papineni et al., 2002), developed from natural language processing, assesses code quality using N-gram precision and a brevity penalty. However, Bleu performs poorly in evaluating code semantics. METEOR (Banerjee and Lavie, 2005) addresses this shortcoming by incorporating external knowledge sources, while CodeBLEU (Ren et al., 2020) improves alignment with human judgment by adding dataflow and AST analysis. Additionally, the application of UML to code summarization (Wang et al., 2021a) introduces CIDEr (Vedantam et al., 2015), a metric from computer vision that combines BLEU with vector space modeling to assess whether the model captures crucial information from UML. Besides, metrics such as BM25 score (Zhang et al., 2023d) and USE (Bansal et al., 2023b, c, a) also utilize algorithms or models to judge similarity.
Model-related metrics contain perplexity, Rlmp, and area under the curve (AUC). Perplexity (Lutellier et al., 2020; Chen et al., 2021a; Lu et al., 2022) quantifies how well a probability model predicts a sample by measuring the average uncertainty or “surprise of the model when encountering new data. Rlmp (Zhang et al., 2023c; Li et al., 2024b) evaluates the improvement of a proposed model compared with baseline models. AUC (Lin et al., 2022; Chen et al., 2023; Xu et al., 2021; Jiang et al., 2024), a widely used metric for evaluating the performance of binary classification models, measures the model’s ability to distinguish between positive and negative classes by calculating the area under the receiver operating characteristic (ROC) curve.
Besides these types of metrics, compiler-based metrics utilize compilers to analyze the quality of predicted code, such as passing compilation process (Agrawal et al., 2024; Wang et al., 2022) or passing test cases (Zan et al., 2023, 2022a; Tang et al., 2023; Zhou et al., 2023; Zhang et al., 2023d; Bi et al., 2024b; Yu et al., 2024). Frequency-based metrics measure how often certain elements or patterns occur, such as normalized inverse word frequency (Zhang et al., 2021) or unchanged identifiers ratio (Jiang et al., 2007). Coverage-based metrics evaluate how thoroughly a model considers relevant conditions, such as inspecting necessary statements (Motwani and Brun, 2023), or how well the generated outputs cover expected elements, such as including libraries (Liao et al., 2024), APIs (Zhang et al., 2024f), or dependencies (Wang et al., 2024g). Higher coverage typically indicates a more comprehensive evaluation. Task-specific metrics craft specific metrics to fit corresponding code intelligence tasks, such as full repair metric (Li et al., 2023c) or fix location filtering score (Kim et al., 2019).
However, due to the lack of suitable metrics or requiring humans to evaluate further, studies (McBurney and McMillan, 2016; Shen et al., 2016; Zhang et al., 2013; Huang et al., 2017; Cortés-Coy et al., 2014; Wong et al., 2015; Chen et al., 2024b) use human evaluation to assess the effectiveness of their methods. For instance, to evaluate the effectiveness of notifications for continuous clone monitoring, study (Zhang et al., 2013) collects feedback from 9 participants to determine if the notifications meet their requirements. Li et al. (Li et al., 2024a) enroll participants to evaluate the quality of commit messages from four dimensions: rationality, comprehensiveness, conciseness, and expressiveness. Developing automated metrics that can better fit relevant studies may enhance the evaluation process and provide more consistent and scalable assessments.
Furthermore, current evaluation metrics predominantly focus on end-to-end performance while paying insufficient attention to context utilization. Neglecting to assess the intermediate steps of context utilization renders these methods a “black box”. This oversight can obscure potential bottlenecks or errors, ultimately complicating the optimization process.

### v2

@@ 7.1. Metric
The evaluation metric is crucial for measuring prediction accuracy and assessing the model’s performance. Table LABEL:tab:Metric summarizes the metrics utilized in all 146 papers and details their respective usage scenarios. We classify metrics based on their function to maintain the table’s clarity and focus. Besides, we exclude some simple evaluation metrics, such as memory usage (Yuan et al., 2020) and time cost (Kamiya, 2015).
@@
Ranking metrics assess models’ effectiveness in recommending solutions for code intelligence tasks. The top@k metric (Zhang et al., 2023d; Bangcharoensap et al., 2012; Yang et al., 2024; Wen et al., 2016; Zhao et al., 2024) measures the proportion of relevant results among the top k recommendations, while MAP (Saha et al., 2013; Lam et al., 2015; Swe and Oo, 2019; Zhou et al., 2012; Wang and Lo, 2014), MAR (Zhang et al., 2023d; Youm et al., 2015), MFR (Zhang et al., 2023d), and MRR (Asaduzzaman et al., 2016; Saha et al., 2013; Lam et al., 2015; Wen et al., 2016; Zhao et al., 2024) offer a more detailed evaluation of the ranking quality.
@@
Classification metrics measure model performance using confusion matrices, which categorize predictions into positive and negative as one dimension and the correctness of the model prediction as another. Accuracy (Zhou et al., 2023; Abhinav et al., 2021; Lam et al., 2015; Zou et al., 2020; Liu et al., 2013), precision (Nafi et al., 2019, 2018; Li et al., 2024b; Ma et al., 2024b; Zhang et al., 2024d; Jimenez et al., 2024), recall (Qiu et al., 2024; Liang et al., 2024; Liao et al., 2024; Xie et al., 2024; Li et al., 2024c) and F1 score (Nguyen et al., 2012; Shahbazi and Fard, 2023; Zou et al., 2020; Liu et al., 2013; Zhou et al., 2019) are four common classification metrics derived from confusion matrices. Matthews correlation coefficient (MCC) (Chen et al., 2023; Qiu et al., 2024; Wen et al., 2024; Jiang et al., 2024) further considers imbalanced datasets and utilizes confusion matrices to summarize classification quality.
Similarity metrics assess how closely the model’s predictions match the ground truth. Bleu (Papineni et al., 2002), developed from natural language processing, assesses code quality using N-gram precision and a brevity penalty. However, Bleu performs poorly in evaluating code semantics. METEOR (Banerjee and Lavie, 2005) addresses this shortcoming by incorporating external knowledge sources, while CodeBLEU (Ren et al., 2020) improves alignment with human judgment by adding dataflow and AST analysis. Additionally, the application of UML to code summarization (Wang et al., 2021b) introduces CIDEr (Vedantam et al., 2015), a metric from computer vision that combines BLEU with vector space modeling to assess whether the model captures crucial information from UML. Besides, metrics such as BM25 score (Zhang et al., 2023c) and USE (Bansal et al., 2023b, c, a) also utilize algorithms or models to judge similarity.
Model-related metrics contain perplexity, Rlmp, and area under the curve (AUC). Perplexity (Lutellier et al., 2020; Chen et al., 2021b; Lu et al., 2022) quantifies how well a probability model predicts a sample by measuring the average uncertainty or “surprise of the model when encountering new data. Rlmp (Zhang et al., 2023d; Li et al., 2024a) evaluates the improvement of a proposed model compared with baseline models. AUC (Lin et al., 2022; Chen et al., 2023; Xu et al., 2021; Jiang et al., 2024), a widely used metric for evaluating the performance of binary classification models, measures the model’s ability to distinguish between positive and negative classes by calculating the area under the receiver operating characteristic (ROC) curve.
Besides, compiler-based metrics use compilers to assess the quality of predicted code, such as by checking whether the code compiles (Agrawal et al., 2024; Wang et al., 2022) or passes test cases (Zan et al., 2023, 2022; Tang et al., 2023; Zhou et al., 2023; Zhang et al., 2023c; Bi et al., 2024; Yu et al., 2024). Frequency-based metrics measure how often certain elements or patterns occur, such as normalized inverse word frequency (Zhang et al., 2021) or unchanged identifiers ratio (Jiang et al., 2007). Coverage-based metrics evaluate how thoroughly a model considers relevant conditions, such as inspecting necessary statements (Motwani and Brun, 2023), or how well the generated outputs cover expected elements, such as including libraries (Liao et al., 2024), APIs (Zhang et al., 2024c), or dependencies (Wang et al., 2024a). Higher coverage typically indicates a more comprehensive evaluation. Task-specific metrics are designed to fit specific code intelligence tasks, such as the full repair metric (Li et al., 2023b) or the fix location filtering score (Kim et al., 2019).
@@
Ranking Top@K (Hit@K) This metric measures the ability of a model to include relevant items in the top K recommendations. (Zhang et al., 2023d; Bangcharoensap et al., 2012; Youm et al., 2015; Swe and Oo, 2019; Zhou et al., 2012; Gao et al., 2023; Motwani and Brun, 2023; Youm et al., 2017; Liu et al., 2023b; Yang et al., 2024; Wen et al., 2016; Zhao et al., 2024) Defect Detection, Code Generation, Program Repair
MAP (@k) Mean Average Precision (MAP) calculates the average precision for each query for each class. MAP@k calculates the same but within the top k results. (Saha et al., 2013; Lam et al., 2015; Swe and Oo, 2019; Zhou et al., 2012; Wang and Lo, 2014; Youm et al., 2017; Yang et al., 2024; Lu et al., 2022; Wen et al., 2016; Zhao et al., 2024) Defect Detection
MAR Mean Average Rank (MAR) denotes the mean of the average rank of finding all items. (Zhang et al., 2023d; Youm et al., 2015) Defect Detection
MFR Mean First Rank (MFR) is the mean of localizing the first faulty statement’s rank of all faults under multiple faulty scenarios. (Zhang et al., 2023d) Defect Detection
@@
Classification Accuracy Accuracy represents the proportion of correctly classified samples to the total number of samples. (Lin et al., 2022; Nashid et al., 2023; Chen et al., 2021b; Asaduzzaman et al., 2016; Raychev et al., 2016; Agrawal et al., 2024; Li et al., 2023b; Zhou et al., 2023; Abhinav et al., 2021; Lam et al., 2015; Zou et al., 2020; Liu et al., 2013, 2023c, 2022, 2023a; Phan et al., 2024; Qiu et al., 2024; Lyu et al., 2021; Nashid et al., 2024; Liang et al., 2024; Liao et al., 2024; Wang et al., 2024c; Xu et al., 2021; Jiang et al., 2024; Zhang et al., 2024d) Program Repair, Code Completion, Code Generation, Defect Detection
@@
Precision (@k) Precision measures the proportion of correctly predicted positive instances out of all instances predicted as positive. Precision @k calculates the same, but within the top k results. (Agashe et al., 2019; Li, 2020; Lin et al., 2022; Wen et al., 2018; Yu et al., 2023; Asaduzzaman et al., 2016; Nguyen et al., 2012; Liu et al., 2013; Zhou et al., 2019; Yuan et al., 2020; Zügner et al., 2021; Bowman and Huang, 2020; Ghosh and Kuttal, 2018; Zhang et al., 2023b; Chen et al., 2023; Liu et al., 2022, 2023a, 2023b; Qiu et al., 2024; Liang et al., 2024; Liao et al., 2024; Wen et al., 2024; Xu et al., 2021; Lu et al., 2022; Wang et al., 2023; Nafi et al., 2019, 2018; Li et al., 2024b; Ma et al., 2024b; Zhang et al., 2024d; Jimenez et al., 2024) Code Generation, Program Repair, Code Completion, Clone Detection, Defect Detection
@@
Recall (@k) Recall represents the proportion of correctly predicted positive samples among all positive samples. Recall@K measures the same, but within the top k suggestions instead of positive samples. (Agashe et al., 2019; Li, 2020; Lin et al., 2022; Yu et al., 2023; Asaduzzaman et al., 2016; Saha et al., 2013; Nguyen et al., 2012; Wang et al., 2009; Zou et al., 2020; Liu et al., 2013; Zhou et al., 2019; Yuan et al., 2020; Zügner et al., 2021; Bowman and Huang, 2020; Ghosh and Kuttal, 2018; Zhang et al., 2023b; Chen et al., 2023; Liu et al., 2022, 2023a, 2023b; Qiu et al., 2024; Liang et al., 2024; Liao et al., 2024; Xie et al., 2024; Li et al., 2024c; Wen et al., 2024; Bibaev et al., 2022; Xu et al., 2021; Nafi et al., 2019, 2018; Li et al., 2024b; Ma et al., 2024b; Zhang et al., 2024d) Code Generation, Program Repair, Code Completion, Defect Detection, Clone Detection
@@
F1-score F1-score comprehensively considers the model’s ability by measuring the harmonic mean of precision and recall. (Liu et al., 2023b; Li, 2020; Zhou et al., 2023; Lin et al., 2022; Yu et al., 2023; Asaduzzaman et al., 2016; Nguyen et al., 2012; Shahbazi and Fard, 2023; Zou et al., 2020; Liu et al., 2013; Zhou et al., 2019; Yuan et al., 2020; Zügner et al., 2021; Bowman and Huang, 2020; Zhang et al., 2023b; Chen et al., 2023; Liu et al., 2022, 2023a; Ding et al., 2023a; Lyu et al., 2021; Liang et al., 2024; Liao et al., 2024; Cheng et al., 2024; Wen et al., 2024; Xu et al., 2021; Jiang et al., 2024; Nafi et al., 2019, 2018; Li et al., 2024b; Zhang et al., 2024d; Liu et al., 2024a) Code Generation, Code Completion, Code Summarization, Program Repair, Clone Detection, Defect Detection
@@
MCC Matthews correlation coefficient (MCC) evaluates binary classification performance using the confusion matrix. This metric provides a balanced score even when the class sizes are imbalanced. (Chen et al., 2023; Qiu et al., 2024; Wen et al., 2024; Jiang et al., 2024) Defect Detection
@@
Similarity BLEU BLEU computes the similarity between prediction and ground truth sequences by assessing the n-gram overlap. (Agashe et al., 2019; Nashid et al., 2023; Yu et al., 2020; Zhou et al., 2023; Washio and Miyao, 2022; Bansal et al., 2023b; Shahbazi and Fard, 2023; Xu et al., 2022; Nie et al., 2021; Wang et al., 2021a; Vu et al., 2023; Shahbazi et al., 2021; Guo et al., 2023b; Ding et al., 2023b; Haque et al., 2020; Wang et al., 2021b; Bansal et al., 2021, 2023c, 2023a; Matmti and Fard, 2023; Lin et al., 2021; Gao et al., 2023; Liu et al., 2021; Zhang et al., 2021; Lyu et al., 2021; Wang et al., 2024a, b; Mandli et al., 2024; Guo et al., 2024b; Wang et al., 2023; Li et al., 2024d; Eliseeva et al., 2023; Imani et al., 2024; Zhang et al., 2024b; Abdu et al., 2024; Zhou et al., 2024) Code Generation, Code Summarization, Program Repair, Commit Message Generation, Code Completion
@@
CodeBLEU CodeBLEU is a bleu-modified metric tailored for code-related tasks. CodeBLEU evaluates the similarity of code by weighting the combination of n-gram match, data-flow match, and AST match. (Liu et al., 2023b; Phan et al., 2024; Wang et al., 2024a; Tan et al., 2024; Li et al., 2024e) Code Generation, Code Completion
@@
EM Exact Match (EM) measures the proportion of predictions that match the ground truth exactly. (Li et al., 2023b; Zhou et al., 2023; Pei et al., 2023; Ding et al., 2023b; Shrivastava et al., 2023b; Zhang et al., 2023a; Shrivastava et al., 2023a; Guo et al., 2023a; Ding et al., 2023a; Liu et al., 2023b; Agashe et al., 2019; Phan et al., 2024; Wu et al., 2024a; Li et al., 2024f; Nashid et al., 2024; Liang et al., 2024; Wang et al., 2024a; Cheng et al., 2024; Tan et al., 2024; Li et al., 2024e; Wang et al., 2024g; Liu et al., 2024b; Deng et al., 2024; Lu et al., 2022; Eliseeva et al., 2023; Prenner and Robbes, 2024; Liu et al., 2024a) Code Generation, Code Completion, Commit Message Generation
@@
ES Edit Similarity (ES) measures the similarity between two strings, defined as the minimum number of single-character edits. (Pei et al., 2023; Shrivastava et al., 2023b; Zhang et al., 2023a; Guo et al., 2023a; Ding et al., 2023a; Wu et al., 2024a; Li et al., 2024f; Liang et al., 2024; Wang et al., 2024a; Cheng et al., 2024; Tan et al., 2024; Wang et al., 2024g; Liu et al., 2024b; Deng et al., 2024; Lu et al., 2022; Eliseeva et al., 2023; Wang et al., 2022; Liu et al., 2024a) Code Generation, Code Completion, Commit Message Generation
SM Syntax Match (SM) quantifies the match of subtrees within the code.  (Li et al., 2024e) Code Completion
@@
ROUGE ROUGE utilizes the longest common subsequence to measure the matching degree between the generated and reference results. (Yu et al., 2020; Xu et al., 2022; Nie et al., 2021; Wang et al., 2021a; Vu et al., 2023; Shahbazi et al., 2021; Zhou et al., 2019; Guo et al., 2023b; Haque et al., 2020; Wang et al., 2021b; Bansal et al., 2021, 2023a; Matmti and Fard, 2023; Lin et al., 2021; Gao et al., 2023; Liu et al., 2021; Zhang et al., 2021; Lyu et al., 2021; Wang et al., 2024b; Mandli et al., 2024; Wang et al., 2023; Li et al., 2024d; Imani et al., 2024; Abdu et al., 2024; Zhou et al., 2024) Code Generation, Code Summarization, Commit Message Generation, Program Repair
@@
METEOR METEOR computes the harmonic mean between unigram precision and recall. (Bansal et al., 2023b; Shahbazi and Fard, 2023; Xu et al., 2022; Nie et al., 2021; Wang et al., 2021a; Vu et al., 2023; Shahbazi et al., 2021; Zhou et al., 2019; Guo et al., 2023b; Wang et al., 2021b; Bansal et al., 2023c, a; Matmti and Fard, 2023; Gao et al., 2023; Liu et al., 2021; Zhang et al., 2021; Wang et al., 2024b; Mandli et al., 2024; Guo et al., 2024b; Li et al., 2024d; Zhang et al., 2024b; Abdu et al., 2024; Zhou et al., 2024; Zhang et al., 2024d) Code Summarization, Commit Message Generation, Program Repair
@@
CIDEr CIDEr is a consensus-based evaluation metric used in image captioning tasks to measure whether the model captures important messages. (Wang et al., 2021b; Lyu et al., 2021) Code Generation, Code Summarization
BM25 score Best Matching 25 (BM25) score estimates the relevance of texts based on term frequency, inverse document frequency, and document length normalization. (Zhang et al., 2023c) Code Generation
RIBES RIBES estimates models based on rank correlation and evaluates the distance between predictions and references. (Lyu et al., 2021) Code Generation
Model-Related Perplexity Perplexity measures how well a probability distribution predicts a sample. This metric computes the exponentiation of the average negative log-likelihood of the predicted probabilities. (Lutellier et al., 2020; Chen et al., 2021b; Lu et al., 2022) Program Repair, Code Completion
RImp Relative Improvement (RImp) calculates the improvement of solving tasks between the old and new techniques. (Zhang et al., 2023d; Li et al., 2024a) Defect Detection, Program Repair
@@
Compiler-Based Pass@k This metric calculates the proportion of solved questions that pass the unit test. (Zan et al., 2023, 2022; Tang et al., 2023; Zhou et al., 2023; Zhang et al., 2023c; Bi et al., 2024; Yu et al., 2024; Xie et al., 2024; Li et al., 2024c; Wang et al., 2024a; Zhang et al., 2024a; Zheng et al., 2024b; Wu et al., 2024b; Guan and Treude, 2024; Zhuo et al., 2024) Code Generation, Code Completion
@@
ValRate Static Validity Rate (ValRate) calculates the performance of models in reducing dependency errors by evaluating the proportion of generated code that successfully passes a static check for dependency errors. (Wang et al., 2024a) Code Generation
@@
ACR Average Coverage Ratio (ACR) calculates the average ratio of successfully predicted statements in all buggy statements. (Yang et al., 2024) Defect Detection
@@
Dependency Coverage Dependency Coverage means the proportion of dependencies that appear in ground-truth functions while covered by the prediction. (Wang et al., 2024a) Code Generation
API Coverage API Coverage reports the ratio of API coverage in code snippets. (Zhang et al., 2024c) Defect Detection
Task-Specific Full Repair Full Repair evaluates the ability to fix a broken program, including the location of an error statement and the fix of this statement. (Li et al., 2023b) Program Repair
@@
However, due to the lack of suitable metrics or the need for human evaluation, studies (McBurney and McMillan, 2016; Shen et al., 2016; Zhang et al., 2013; Huang et al., 2017; Cortés-Coy et al., 2014; Wong et al., 2015; Chen et al., 2024) use human evaluation to assess the effectiveness of their methods. For instance, to evaluate the effectiveness of notifications for continuous clone monitoring, study (Zhang et al., 2013) collects feedback from 9 participants to determine if the notifications meet their requirements. Li et al. (Li et al., 2024d) enroll participants to evaluate the quality of commit messages from four dimensions: rationality, comprehensiveness, conciseness, and expressiveness. Developing automated metrics that better align with relevant studies may enhance the evaluation process and provide more consistent, scalable assessments.
Furthermore, current evaluation metrics predominantly focus on end-to-end performance while paying insufficient attention to context utilization. Neglecting to assess the intermediate steps of context utilization renders these methods a “black box”. This oversight can obscure potential bottlenecks or errors, ultimately complicating the optimization process. To address this, developing more fine-grained metrics for context utilization is necessary. For example, RepoExec (Hai et al., 2024) proposes “Dependency Invocation Rate” to directly measure how well an LLM utilizes the provided source code context to generate code with dependencies.